it tell you to use the substitution method to solve the linear system
x +2y = - 5
4x-3y=2

3x-2y=4
x+3y=5

3x+y= -2
x+3y=2

Answer by K N Swamy(1)

q1.
x+2y=-5----(1)
4x-3y=2---(2)
from (1), X=-2Y-5
substitute x=-2y-5 in (2)
4(-2y-5)-3y=2
-8y-20-3y=2
-11y-20=2
-11y=-22
y=-22/-11
y=2
x=-2y-5
substitute y=2
x=-2(2)-5
x=-4-5
x=-9
Hence x=-9 and y=2